# A Heterogeneous Dynamic Convolutional Neural Network for Image Super-resolution

### v2

@@ IV-A Training Dataset
We utilize the public DIV2K dataset[38] as a training set to develop a HDSRNet for enhancing image super-resolution capabilities. Specifically, the DIV2K contains three parts, i.e., a training dataset of 800 images, a validation dataset of 100 images and a test dataset of 100 images. To enlarge diversity of a training dataset, we merge an original dataset and a validation dataset to form a new training dataset, which includes 900 high-quality images and corresponding low-resolution images of x2, x3 and x4. All training images are saved format of ‘.png’.
@@ IV-B Test Datasets
To fairly test SR performance of our HDSRNet, four public SR datasets, i.e., Set5 containing 5 natural images [39], Set14 containing 14 natural images [40], BSD100 (B100) containing 100 natural images [41] and Urban100 (U100) containing 100 containing [42] are used to conduct experiments. These datasets include three different scales, i.e., x2, x3 and x4. They are saved as format of ‘.png’.
@@ IV-C Implementation Details
Original parameters of training a HDSRNet are $\beta 1$ of 0.9, $\beta 2$ of 0.999, epsilon of 1e-8, batch size of 64 and initial learning rate of 1e-4, which is halved every 300,000 steps. Also, it more parameters are the same as Ref. [19].
The proposed HDSRNet is implemented by Pytorch 1.8.0 and Python 3.8. And all the experiments run on a workstation with Ubuntu 20.04, which equipped AMD EPYC 7502P CPU and four GPUs of Nvidia GeForce RTX 3090 with Nvidia CUDA 11.1, and cuDNN 8005. All experiments are trained via one 3090 GPU.
@@ IV-D Network Analysis
We analyze architecture of designed HDSRNet containing a heterogeneous upper sub-network, symmetrical lower-network and construction block for image super-resolution, according to its rationality and validity.
@@
Heterogeneous upper sub-network: Most of existing SR models cannot adaptively learn parameters, according to different given low-resolution images [43]. Dynamic convolutions can automatically adjust parameters, according to different inputs [26]. In response to this motivation, we have developed a heterogeneous upper sub-network specifically for image super-resolution. We design a main network, according to VGG [44]. That is, six stacked Conv+ReLU is used as a basic network to extract structural low-frequency information. A combination of six stacked Conv+ReLU and a CB has obtained peak signal-to-noise ratio (PSNR) [45] of 25.13dB and structural similarity (SSIM) [45] of 0.7525 on U100 for x4, which shows effectiveness of six stacked Conv+ReLU. A CB is used to build HR images, which can be shown in latter section. To prevent long-term dependency problem, deep fusion idea [46] is used in this paper. That is, each residual learning operation is used to act between an input and output of Conv+ReLU besides the first Conv+ReLU to transform low-frequency information from shallow layers to deep layers to pursue better performance on super-resolution. Its effectiveness can be shown by comparing ‘Heterogeneous upper sub-network without Dilated Conv+ReLU, Dynamic Conv+ReLU and a CB’ and ‘A combination of six stacked Conv’ in terms of PSNR and SSIM in TABLE I. To extract more context, each dilated Conv+ReLU with dilated factor of 2 is used before each Conv+ReLU besides the first Conv+ReLU in heterogeneous upper sub-network and its effectiveness can be proved by comparing ‘Heterogeneous upper sub-network without Dilated Conv+ReLU, Dynamic Conv+ReLU and a CB’ and ‘Heterogeneous upper sub-network without dynamic Conv+ReLU and a CB’ in terms of PSNR and SSIM on U100 for x4. Specifically, five stacked blocks besides the first Conv+ReLU in the heterogeneous upper sub-network (HUNet) are five heterogeneous blocks (HBs) and its rationality and validity are shown in latter section. To make HUNet robust, each Dynamic Conv+ReLU is set behind each Dilated Conv+ReLU in each HB to dynamically adjust parameters, according to different inputs. Its better SR results can be found via comparing ‘Heterogeneous upper sub-network and a CB’ and ‘Heterogeneous upper sub-network without dynamic Conv+ReLU and a CB’ in TABLE I, where effectiveness of dynamic convolutions is verified. It is known that difference of network architecture is bigger, its performance is better [25]. Motivated by that, we respectively use Conv+ReLU rather than Dynamic Conv+ReLU and Dilated Conv+ReLU to conduct experiments. As shown in TABLE I, we can see that ‘Heterogeneous upper sub-network and a CB’ has obtained higher PSNR and SSIM values than that of ‘Heterogeneous upper sub-network with Conv+ReLU rather than dilated Conv+ReLU’ and ‘Heterogeneous upper sub-network with Conv+ReLU rather than dynamic Conv+ReLU’. This approach demonstrates the effectiveness of dynamic Conv+ReLU and dilated Conv+ReLU in HB, and highlights the advantages of various network architectures within HB for image super-resolution.
@@
TABLE I: PSNR and SSIM of different methods on U100 for x4.
@@
×4 A combination of six stacked Conv+ReLU and a CB 25.13/0.7525
@@
Conv+ReLU, Dynamic Conv+ReLU and a CB
@@
Conv+ReLU and a CB
Heterogeneous upper sub-network with Conv+ReLU 25.74/0.7739
rather than dilated Conv+ReLU
Heterogeneous upper sub-network with Conv+ReLU 25.68/0.7729
rather than dynamic Conv+ReLU
@@
HDSRNet without residual learning operations 25.99/0.7825
@@
HDSRNet 26.01/0.7827
TABLE II: PSNR and SSIM of different methods with different upscale factors on Set5.
@@
HDSRNet (Ours) 37.95/0.9604 34.31/0.9263 32.15/0.8927
TABLE III: PSNR and SSIM of different methods with different upscale factors on Set14.
@@
HDSRNet (Ours) 33.52/0.9170 30.27/0.8412 28.56/0.7799
TABLE IV: PSNR and SSIM of different methods with different upscale factors on B100.
@@
HDSRNet (Ours) 32.14/0.8994 29.06/0.8048 27.55/0.7357
TABLE V: PSNR and SSIM of different methods with different upscale factors on U100.
@@
Image sizes VDSR [13] CARN-M [23] ACNet [66] HDSRNet (Ours)
@@
HDSRNet (Ours) 1,819K 110.99G
Figure 2: Predicted high-quality images from different methods on a same low-resolution image from Set14 for ×2: (a) A HR image, (b)Bicubic, (c) SRCNN, (d) LESRCNN, (e) DCLS, (f) VDSR, (g) DRCN and (h) HDSRNet (Ours).
Figure 3: Predicted high-quality images from different methods on a same low-resolution image from B100 for ×3: (a) A HR image, (b)Bicubic, (c) SRCNN, (d) LESRCNN, (e) DCLS, (f) VDSR, (g) DRCN and (h) HDSRNet (Ours).
Figure 4: Predicted high-quality images from different methods on a same low-resolution image from B100 for ×3: (a) A HR image, (b)Bicubic, (c) SRCNN, (d) LESRCNN, (e) DCLS, (f) VDSR, (g) DRCN and (h) HDSRNet (Ours).
Figure 5: Predicted high-quality images from different methods on a same low-resolution image from U100 for ×4: (a) A HR image, (b)Bicubic, (c) SRCNN, (d) LESRCNN, (e) DCLS, (f) VDSR, (g) DRCN and (h) HDSRNet (Ours).
Figure 6: Predicted high-quality images from different methods on a same low-resolution image from U100 for ×4: (a) A HR image, (b)Bicubic, (c) SRCNN, (d) LESRCNN, (e) DCLS, (f) VDSR, (g) DRCN and (h) HDSRNet (Ours).
Figure 7: Predicted high-quality images from different methods on a same low-resolution image from U100 for ×4: (a) A HR image, (b)Bicubic, (c) SRCNN, (d) LESRCNN, (e) DCLS, (f) VDSR, (g) DRCN and (h) HDSRNet (Ours).
The first phase design a 16-layer stacked Conv+ReLU to extract low-frequency structural information, according to VGG [44]. Its effectiveness can be proved via ‘Heterogeneous upper sub-network and a CB’ and ‘HDSRNet without residual learning operations in symmetrical lower sub-network’ in TABLE I. To enhance relationship of hierarchical features, the second phase is conducted. It uses residual learning operations to merge obtained information of shallow and deep layers to achieve a symmetrical lower sub-network to extract more accurate information as shown in Section III.D. As illustrated in TABLE I, we can see that ‘HDSRNet’ has obtained higher PSNR and SSIM than that of HDSRNet without residual learning operations in symmetrical lower sub-network, which shows effectiveness of residual learning operations in the symmetrical lower sub-network for image SR. Besides, ’HDSRNet’ has obtained higher PSNR value than that of ’Heterogeneous upper sub-network and a CB’ in TABLE I, which shows effectiveness of Heterogeneous parallel networks for image super-resolution.
Construction block: To construct HR images, we design a 2-layer construction block. It consists of a sub-pixel convolutional layer and a convolutional layer. Sub-pixel convolutional layer is used to amplify low-frequency information to high-frequency information. A convolutional layer is utilized to construct predicted HR images.
@@ IV-E Comparisons with Popular Methods for SR
In this section, we use both quantitative and qualitative analysis to evaluate results of our HDSRNet on SISR. Quantitative analysis includes PSNR, SSIM, running time and complexity of our HDSRNet. Specifically, PSNR and SSIM are used to measure quality of predicted HR images. Also, running time and parameters are used to test feasibility of our HDSRNet for real applications, i.e., phones and cameras. We use A+ [47], jointly optimized regressors (JOR) [48], image upscaling with super-resolution forests (RFL) [49], self-exemplars SR method (SelfEx) [42], cascade of sparse coding based network (CSCN) [50], image restoration using encoder-decoder networks (RED) [51], denoising convolutional neural network (DnCNN) [52], trainable non-linear reaction diffusion (TNRD) [53], fast dilated residual SR convolutional net-work (FDSR) [54], SRCNN [12], FSRCNN [18], residue context sub-network (RCN) [55], VDSR [13], deeply-recursive convolutional network (DRCN) [14], IDN [34], DRRN [15], Laplacian SR network (LapSRN) [56], new architecture of deep recursive convolution networks for SR (NDRCN) [57], a persistent memory network (MemNet) [22], CARN-M [23], light-weight image super-resolution with enhanced CNN (LESRCNN) [58], deep alternating network (DAN) [61], Pixel-Level Generative Adversarial Network (PGAN) [62] and our HDSRNet on four public datasets, i.e., Set5, Set14, B100 and U100 for x2, x3 and x4 to conduct experiments. As shown in TABLEs II and III, we can see that our HDSRNet has obtained the best performance in terms of PSNR and SSIM for x2, x3 and x4. For instance, our HDSRNet has improvements of 0.12dB on PSNR and 0.004 on SSIM on Set 5 for x2 than that of IDN in TABLE II. Our HDSRNet has obtained improvements of 0.1dB on PSNR and 0.003 on SSIM on Set14 for x4 than that of DSRCNN in TABLE III. That also shows that our HDSRNet is effective on small datasets for image super-resolution. For big datasets, our HDSRNet still has an advantage for image super-resolution in TABLEs IV and V, we can see that our HDSRNet has obtained the best SR results for all the scale factors, i.e., x2, x3 and x4. For instance, our HDSRNet has exceeded 0.09dB on PSNR and 0.0016 on SSIM than that of DSRCNN for x2 on B100 in TABLE IV. Our HDSRNet has exceeded 0.07dB on PSNR and 0.0012 on SSIM than that of DSRCNN for x4 on U100 in TABLE V. That shows that our HDSRNet is suitable to big datasets for image super-resolution. Specifically, red and blue lines denote the best and second SR results from TABLE II to TABLE V, respectively.
@@
To test practicality of our HDSRNet, we use running time and complexity to test performance of our HDSRNet for image SR. As shown in TABLE VI, our HDSRNet has obtained competitive running time for restoring low-resolution images sizes of $256\times 256$ and $512\times 512$. As illustrated in TABLE VII, although our HDSRNet has obtained more parameters than that of ACNet, it has obtained less flops than that of ACNet. Thus, it is competitive in complexity and suitable for application in consumer electronic products. According to mentioned experiment results, we can find that our HDSRNet is effective in terms of quantitative analysis.
@@
For qualitative analysis, we use Bicubic, SRCNN, LESRCNN, DCLS, VDSR, DRCN and HDSRNet on a low-resolution image from the B100 and U100 for x3 and x4 to recover high-quality images, which are used to compare with given HR images. That is, we amplify one area of predicted high-quality images from different methods as observation areas, observation areas are clearer, their corresponding methods are more effective for image super-resolution. As shown in Figs.2-7, we can see that our HDSRNet has obtained clearer areas, it shows that our HDSRNet is effective for qualitative analysis. In a summary, our HDSRNet is a good tool for image resolution, according to quantitative and qualitative analysis.

### v3

@@ IV-A Training Dataset
We utilize the public DIV2K dataset[40] as a training set to develop a ADSRNet for enhancing image super-resolution capabilities. Specifically, the DIV2K contains three parts, i.e., a training dataset containing 800 images, a validation dataset containing 100 images and a test dataset containing 100 images. To enlarge diversity of a training dataset, we merge an original dataset and a validation dataset to form a new training dataset, which includes 900 high-quality images and corresponding LR images of x2, x3 and x4. All training images are saved format of ‘.png’.
@@ IV-B Test Datasets
To fairly test SR performance of our ADSRNet, four public SR datasets, i.e., Set5 containing 5 natural images [41], Set14 containing 14 natural images [42], BSD100 (B100) containing 100 natural images [43] and Urban100 (U100) containing 100 containing [44] are used to conduct experiments. These datasets include three different scales, i.e., x2, x3 and x4. They are saved as format of ‘.png’.
@@ IV-C Implementation Details
Original parameters of training a ADSRNet are $\beta 1$ of 0.9, $\beta 2$ of 0.999, epsilon of 1e-8, batch size of 64 and initial learning rate of 1e-4, which is halved every 300,000 steps. Also, it more parameters are the same as Ref. [22].
The proposed ADSRNet is developed based on Pytorch 1.8.0 and Python 3.8. And all the experiments run on a workstation with Ubuntu 20.04, which equipped AMD EPYC 7502P CPU and four GPUs of Nvidia GeForce RTX 3090 with Nvidia CUDA 11.1, and cuDNN 8005. All experiments are trained via one 3090 GPU.
@@ IV-D Network Analysis
We analyze architecture of designed ADSRNet containing a heterogeneous upper sub-network, symmetrical lower-network and construction block for image super-resolution, according to its rationality and validity.
@@
Heterogeneous upper sub-network: Most of existing SR models cannot adaptively learn parameters, according to different given low-resolution images [45]. Dynamic convolutions can automatically adjust parameters, according to different inputs [29]. In response to this motivation, we have developed a heterogeneous upper sub-network specifically for image super-resolution. A heterogeneous network architecture can facilitate more diversified structural information, which can provide complementary information to strengthen robustness of an obtained SISR model. Efficient super-resolution performance of proposed ADSRNet is implemented by enhancing relation of context information, salient information relation of a kernel mapping and relations of shallow and deep layers. That is, dilated convolutions are used to enhance relations of context information for SISR. Dynamic convolutions are used to enhance relations of salient information in a kernel mapping to improve adaptive ability of obtained ADSRNet. Also, residual learning technique is used to enhance relations of shallow and deep layers to prevent long-term dependency issue for SISR. More detailed design of the heterogeneous upper sub-network is given as follows. We design a main network, according to VGG [46]. That is, six stacked CRU is used as a basic network to extract structural low-frequency information. A combination of six stacked CRU and a CB has obtained peak signal-to-noise ratio (PSNR) [47] of 25.13dB and structural similarity (SSIM) [47] of 0.7525 on U100 for x4, which shows effectiveness of six stacked CRU. A CB is used to build HR images, which can be shown in latter section. To prevent long-term dependency problem, deep fusion idea [48] is used in this paper. That is, each RL operation is used to act between an input and output of CRU besides the first CRU to transform low-frequency information from shallow layers to deep layers to pursue better performance on super-resolution. Its effectiveness can be shown by comparing ‘Heterogeneous upper sub-network without Dilated CRU, Dynamic CRU and a CB’ and ‘A combination of six stacked Conv’ in terms of PSNR and SSIM in TABLE I. To extract more context, each dilated CRU with dilated factor of 2 is used before each CRU besides the first CRU in heterogeneous upper sub-network. Its effectiveness can be proved by comparing ‘Heterogeneous upper sub-network without Dilated CRU, Dynamic CRU and a CB’ and ‘Heterogeneous upper sub-network without dynamic CRU and a CB’ in terms of PSNR and SSIM on U100 for x4. Specifically, five stacked blocks besides the first CRU in the heterogeneous upper sub-network (HUNet) are five heterogeneous blocks (HBs) rationality and validity are shown in latter section. To make HUNet robust, each Dynamic CRU is set behind each Dilated CRU in each HB to dynamically adjust parameters, according to different inputs. Its better SR results can be found via comparing ‘Heterogeneous upper sub-network and a CB’ and ‘Heterogeneous upper sub-network without dynamic CRU and a CB’ in TABLE I, where effectiveness of dynamic convolutions is verified. It is known that difference of network architecture is bigger, its performance is better [28]. Motivated by that, we respectively use CRU rather than Dynamic CRU and Dilated CRU to conduct experiments. TABLE I shows that ‘Heterogeneous upper sub-network and a CB’ has obtained higher PSNR and SSIM values than that of ‘Heterogeneous upper sub-network with CRU rather than dilated CRU’ and ‘Heterogeneous upper sub-network with CRU rather than dynamic CRU’. This approach demonstrates the effectiveness of dynamic CRU and dilated CRU in HB, and highlights the advantages of various network architectures within HB for image super-resolution.
@@
TABLE I: SISR results of different methods on U100 for x4.
@@
×4 A combination of six stacked CRU and a CB 25.13/0.7525
@@
CRU, Dynamic CRU and a CB
@@
CRU and a CB
Heterogeneous upper sub-network with CRU 25.74/0.7739
rather than dilated CRU
Heterogeneous upper sub-network with CRU 25.68/0.7729
rather than dynamic CRU
@@
ADSRNet without residual learning operations 25.99/0.7825
@@
ADSRNet 26.01/0.7827
TABLE II: SISR results of different methods with on Set5 for three upscale factors.
@@
ADSRNet (Ours) 37.95/0.9604 34.31/0.9263 32.15/0.8927
TABLE III: SISR results of different methods on Set14 for three upscale factors.
@@
ADSRNet (Ours) 33.52/0.9170 30.27/0.8412 28.56/0.7799
TABLE IV: SISR results of different methods on B100 for three upscale factors.
@@
ADSRNet (Ours) 32.14/0.8994 29.06/0.8048 27.55/0.7357
TABLE V: SISR results of different methods on U100 for three upscale factors.
@@
Image sizes VDSR [16] CARN-M [26] ACNet [68] ADSRNet (Ours)
@@
ADSRNet (Ours) 1,819K 110.99G
Figure 2: Predicted high-quality images from different methods on a same low-resolution image from Set14 for ×2: (a) A HR image, (b)Bicubic, (c) SRCNN, (d) LESRCNN, (e) DCLS, (f) VDSR, (g) DRCN and (h) ADSRNet (Ours).
Figure 3: Predicted high-quality images from different methods on a same low-resolution image from B100 for ×3: (a) A HR image, (b)Bicubic, (c) SRCNN, (d) LESRCNN, (e) DCLS, (f) VDSR, (g) DRCN and (h) ADSRNet (Ours).
Figure 4: Predicted high-quality images from different methods on a same low-resolution image from B100 for ×3: (a) A HR image, (b)Bicubic, (c) SRCNN, (d) LESRCNN, (e) DCLS, (f) VDSR, (g) DRCN and (h) ADSRNet (Ours).
Figure 5: Predicted high-quality images from different methods on a same low-resolution image from U100 for ×4: (a) A HR image, (b)Bicubic, (c) SRCNN, (d) LESRCNN, (e) DCLS, (f) VDSR, (g) DRCN and (h) ADSRNet (Ours).
Figure 6: Predicted high-quality images from different methods on a same low-resolution image from U100 for ×4: (a) A HR image, (b)Bicubic, (c) SRCNN, (d) LESRCNN, (e) DCLS, (f) VDSR, (g) DRCN and (h) ADSRNet (Ours).
Figure 7: Predicted high-quality images from different methods on a same low-resolution image from U100 for ×4: (a) A HR image, (b)Bicubic, (c) SRCNN, (d) LESRCNN, (e) DCLS, (f) VDSR, (g) DRCN and (h) ADSRNet (Ours).
The first phase design a 16-layer stacked CRU to extract low-frequency structural information, according to VGG [46]. Its effectiveness can be proved via ‘Heterogeneous upper sub-network and a CB’ and ‘ADSRNet without RL operations in symmetrical lower sub-network’ in TABLE I. To enhance relationship of hierarchical features, the second phase is conducted. It uses RL operations to gather obtained features of shallow and deep layers to achieve a symmetrical lower sub-network to extract more accurate information as shown in Section III.D. As illustrated in TABLE I, we can see that ‘ADSRNet’ has obtained higher PSNR and SSIM than that of ADSRNet without RL operations in symmetrical lower sub-network, which shows effectiveness of RL operations in the symmetrical lower sub-network for image SR. Besides, ’ADSRNet’ has obtained higher PSNR value than that of ’Heterogeneous upper sub-network and a CB’ in TABLE I, which shows effectiveness of Heterogeneous parallel networks for image super-resolution.
Construction block: To construct HR images, we design a 2-layer construction block. It consists of a sub-pixel convolutional layer and a convolutional layer. Sub-pixel convolutional layer is used to amplify low-frequency features to high-frequency features. A convolutional layer is utilized to construct predicted HR images.
@@ IV-E Comparisons with Popular Methods for SR
A heterogeneous network architecture can more diversified structural information, which can provide complementary information to strengthen robustness of an obtained SISR model. The heterogeneous network is deigned, according to enhancing relation of context information, salient information relation of a kernel mapping and relations of shallow and deep layers. Besides, a deep network via parallel networks can extract extra structural information to facilitate more representative information to further improve SR effect. These make our ADSRNet obtain better SR results than these of some classical SR methods, i.e., VDSR [16], DRCN [17]. Rationality and effectiveness of key techniques in our ADSRNet are given in Section IV. D. Superiority of the proposed ADSRNet for SISR is given as follows. In this section, we use both quantitative and qualitative analysis to evaluate results of our ADSRNet on SISR. Quantitative analysis includes PSNR, SSIM, running time and complexity of our ADSRNet. Specifically, PSNR and SSIM are used to measure quality of predicted HR images. Also, running time and parameters are used to test feasibility of our ADSRNet for real applications, i.e., phones and cameras. We use A+ [49], jointly optimized regressors (JOR) [50], image upscaling with super-resolution forests (RFL) [51], self-exemplars SR method (SelfEx) [44], cascade of sparse coding based network (CSCN) [52], image restoration using encoder-decoder networks (RED) [53], denoising convolutional neural network (DnCNN) [54], trainable non-linear reaction diffusion (TNRD) [55], fast dilated residual SR convolutional net-work (FDSR) [56], SRCNN [15], FSRCNN [21], residue context sub-network (RCN) [57], VDSR [16], DRCN [17], IDN [36], DRRN [18], Laplacian SR network (LapSRN) [58], new architecture of deep recursive convolution networks for SR (NDRCN) [59], a persistent memory network (MemNet) [25], CARN-M [26], light-weight image super-resolution with enhanced CNN (LESRCNN) [60], deep alternating network (DAN) [63], Pixel-Level Generative Adversarial Network (PGAN) [64] and our ADSRNet on four public datasets, i.e., Set5, Set14, B100 and U100 for x2, x3 and x4 to conduct experiments. As shown in TABLEs II and III, we can see that our ADSRNet has obtained the best performance in terms of PSNR and SSIM for x2, x3 and x4. For instance, our ADSRNet has improvements of 0.12dB on PSNR and 0.004 on SSIM on Set 5 for x2 than that of IDN in TABLE II. Our ADSRNet has obtained improvements of 0.1dB on PSNR and 0.003 on SSIM on Set14 for x4 than that of DSRCNN in TABLE III. That also shows that our ADSRNet is effective on small datasets for image super-resolution. For big datasets, our ADSRNet still has an advantage for image super-resolution in TABLEs IV and V, we can see that our ADSRNet has obtained the best SR results for all the scale factors, i.e., x2, x3 and x4. For instance, our ADSRNet has exceeded 0.09dB on PSNR and 0.0016 on SSIM than that of DSRCNN for x2 on B100 in TABLE IV. Our ADSRNet has exceeded 0.07dB on PSNR and 0.0012 on SSIM than that of DSRCNN for x4 on U100 in TABLE V. That shows that our ADSRNet is suitable to big datasets for image super-resolution. Specifically, red and blue lines denote the best and second SR results from TABLE II to TABLE V, respectively.
@@
To test practicality of our ADSRNet, we use running time and complexity to test performance of our ADSRNet for image SR. As shown in TABLE VI, our ADSRNet has obtained competitive running time for restoring LR images sizes of $256\times 256$ and $512\times 512$. As illustrated in TABLE VII, although our ADSRNet has obtained more parameters than that of ACNet, it has obtained less flops than that of ACNet. Thus, it is competitive in complexity and suitable for application in consumer electronic products. According to experiment results, our ADSRNet is effective in terms of quantitative analysis.
@@
For qualitative analysis, we use Bicubic, SRCNN, LESRCNN, DCLS, VDSR, DRCN and ADSRNet on a low-resolution image from the B100 and U100 for x3 and x4 to recover high-quality images, which are used to compare with given HR images. That is, we amplify one area of predicted high-quality images from different methods as observation areas, observation areas are clearer, their corresponding methods are more effective for SISR. As shown in Figs.2-7, our ADSRNet has obtained clearer areas, it shows that our ADSRNet is effective for qualitative analysis. In a summary, our ADSRNet is a good tool for image resolution, according to quantitative and qualitative analysis.